# Cardinality-Preserving Attention Channels for Graph Transformers in Molecular Property Prediction

### v3

@@ 3.5 Chemistry-Aware Augmentation Ablations
Attribute masking 0.818 $\pm$ 0.007 0.803 $\pm$ 0.008 +0.015 0.542 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.048 0.920 $\pm$ 0.008
Valency-constrained 0.820 $\pm$ 0.007 0.804 $\pm$ 0.008 +0.016 0.540 $\pm$ 0.02 0.590 $\pm$ 0.03 -0.050 0.922 $\pm$ 0.008
@@ 3.7 Pretraining Corpus and Standardization
  • Source datasets: Random sample of 40M from ZINC20 clean-leads (total 220M SMILES) [8] and full ChEMBL35 (2.4M molecules) [19].
@@ 3.8 Optimization and Training Schedule
Per-view SPD recomputation (via multi-source BFS) increases pretraining cost by  15% in wall-clock time; ablations under reduced budgets (10%, 25%, 100%) show CPA gains persist (Table 13).
@@ 3.9 Faithful Graphormer Baseline
For faithful comparison, we implement Graphormer with global attention ($K=\infty$), full SPD binning (up to max observed, clipped at 20, covering ¿99% SPDs in our corpus), shortest-path edge encodings as learnable biases for each edge type along the shortest path (see appendix for the exact path-bias implementation), and centrality biases. Pretraining uses the same dual objectives (mask + contrast) and corpus. Hyperparameters match our backbone (12 layers, 512 dim, 8 heads). This isolates sparsity and CPA effects from bias design. See appendix for full implementation details. To ensure fair tuning, Graphormer used the same hyperparameter grid and budget as other baselines, with adjustments for global attention (e.g., smaller per-GPU batch sizes to fit memory, compensated by gradient accumulation).
@@ 4 Experiments
All models share architecture, hyperparameters, pretraining schedule, corpus, and fine-tuning protocol. Differences are summarized in Table 5.
@@ 4.3 Efficiency
Table 6: Efficiency measured on A100 80GB GPUs (4 GPUs, DP=data parallel, per-GPU batch=1024, global batch=4096), bf16 (bfloat16 mixed precision), no activation checkpointing, FlashAttention-2 kernel. Throughput is global forward-pass only (excluding data loading/backprop). Peak memory is per-GPU CUDA allocated. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling in the padded attention length $L$ (Table 7). For typical drug-like molecules, $K=3$ is near-global (median coverage ¿95%); substantial sparsity benefits emerge on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 6 reports measured efficiency. Gains are meaningful on large molecules; on small/medium (dominant in benchmarks), sparsity is limited. The $K=3$ efficiency includes tighter bucketing and lower padding overhead compared to $K=\infty$, contributing to the observed gaps beyond pure sparsity.
@@ 4.4 Results
Statistical significance uses paired bootstrap (10,000 resamples per task/seed, with identical resample indices across models to preserve pairing, deltas averaged across seeds); denotes 95% CI excludes zero vs SparseGraphormer-K3. Holm correction is applied across the 11 benchmark tasks for the primary CPA-vs-noCPA comparisons (conclusions unchanged).
@@
Task (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ESOL (RMSE $\downarrow$) 0.598 $\pm$ 0.03 -0.056 0.580 $\pm$ 0.02 0.612 $\pm$ 0.03 0.612 $\pm$ 0.03 0.550 $\pm$ 0.02 0.589 $\pm$ 0.03 0.542 $\pm$ 0.02
Lipophilicity (RMSE $\downarrow$) 0.658 $\pm$ 0.02 -0.029 0.662 $\pm$ 0.02 0.672 $\pm$ 0.02 0.672 $\pm$ 0.02 0.640 $\pm$ 0.01 0.650 $\pm$ 0.02 0.629 $\pm$ 0.01
BBBP (ROC-AUC $\uparrow$) 0.918 $\pm$ 0.011 +0.020 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.910 $\pm$ 0.012 0.930 $\pm$ 0.009 0.915 $\pm$ 0.011 0.938 $\pm$ 0.009
Tox21 (ROC-AUC $\uparrow$) 0.908 $\pm$ 0.009 +0.013 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.902 $\pm$ 0.010 0.915 $\pm$ 0.008 0.905 $\pm$ 0.009 0.921 $\pm$ 0.008
ClinTox (ROC-AUC $\uparrow$) 0.939 $\pm$ 0.017 +0.015 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.932 $\pm$ 0.018 0.952 $\pm$ 0.010 0.940 $\pm$ 0.016 0.954 $\pm$ 0.011
@@
Table 8: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
@@
Dataset (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
ogbg-molhiv (ROC-AUC $\uparrow$) 0.802 $\pm$ 0.009 +0.017 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.800 $\pm$ 0.010 0.810 $\pm$ 0.007 0.795 $\pm$ 0.009 0.819 $\pm$ 0.007
ogbg-molpcba (AP $\uparrow$) 0.294 $\pm$ 0.002 +0.010 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.300 $\pm$ 0.003 0.302 $\pm$ 0.002 0.290 $\pm$ 0.002 0.304 $\pm$ 0.002
@@
Table 9: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks). Additional baselines reproduced under our protocol.
@@
Benchmark (metric) SparseGraphormer-K3 $\Delta$ GraphMAE-K3 GraphCL Graphormer D-MPNN GIN CardinalGraphFormer
Caco2_Wang (MAE $\downarrow$) 0.265 $\pm$ 0.009 -0.024 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.270 $\pm$ 0.009 0.250 $\pm$ 0.007 0.260 $\pm$ 0.008 0.241 $\pm$ 0.007
hERG (ROC-AUC $\uparrow$) 0.870 $\pm$ 0.005 +0.028 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.865 $\pm$ 0.006 0.880 $\pm$ 0.004 0.868 $\pm$ 0.005 0.898 $\pm$ 0.004
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.525 $\pm$ 0.020 +0.013 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.520 $\pm$ 0.021 0.530 $\pm$ 0.017 0.522 $\pm$ 0.019 0.538 $\pm$ 0.017
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.614 $\pm$ 0.012 +0.007 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.610 $\pm$ 0.013 0.618 $\pm$ 0.011 0.612 $\pm$ 0.012 0.621 $\pm$ 0.011
@@
Table 10: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3. Additional baselines reproduced under our protocol.
@@ 4.5 Size-Controlled and Size-Shift Evaluations
ESOL (RMSE $\downarrow$) Small 0.550 $\pm$ 0.02 0.610 $\pm$ 0.03 0.850 $\pm$ 0.05
Mid 0.540 $\pm$ 0.02 0.600 $\pm$ 0.03 0.830 $\pm$ 0.05
Large 0.530 $\pm$ 0.02 0.585 $\pm$ 0.03 0.820 $\pm$ 0.05
Size-shift 0.545 $\pm$ 0.02 0.605 $\pm$ 0.03 0.840 $\pm$ 0.05
BBBP (AUC $\uparrow$) Small 0.935 $\pm$ 0.009 0.915 $\pm$ 0.011 0.650 $\pm$ 0.02
Mid 0.937 $\pm$ 0.009 0.917 $\pm$ 0.011 0.630 $\pm$ 0.02
Large 0.940 $\pm$ 0.009 0.920 $\pm$ 0.011 0.620 $\pm$ 0.02
Size-shift 0.938 $\pm$ 0.009 0.918 $\pm$ 0.011 0.640 $\pm$ 0.02
hERG (AUC $\uparrow$) Small 0.895 $\pm$ 0.004 0.865 $\pm$ 0.005 0.700 $\pm$ 0.01
Mid 0.897 $\pm$ 0.004 0.868 $\pm$ 0.005 0.685 $\pm$ 0.01
Large 0.900 $\pm$ 0.004 0.875 $\pm$ 0.005 0.680 $\pm$ 0.01
Size-shift 0.898 $\pm$ 0.004 0.870 $\pm$ 0.005 0.690 $\pm$ 0.01
Tox21 (AUC $\uparrow$) Small 0.918 $\pm$ 0.008 0.905 $\pm$ 0.009 0.750 $\pm$ 0.01
Mid 0.920 $\pm$ 0.008 0.907 $\pm$ 0.009 0.740 $\pm$ 0.01
Large 0.925 $\pm$ 0.008 0.910 $\pm$ 0.009 0.730 $\pm$ 0.01
Size-shift 0.921 $\pm$ 0.008 0.908 $\pm$ 0.009 0.740 $\pm$ 0.01
molpcba (AP $\uparrow$) Small 0.302 $\pm$ 0.002 0.292 $\pm$ 0.002 0.200 $\pm$ 0.01
Mid 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
Large 0.306 $\pm$ 0.002 0.296 $\pm$ 0.002 0.180 $\pm$ 0.01
Size-shift 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.01
@@ 4.6.1 Cardinality-Control Baselines
Per-view SPD (new) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Stable SPD (old) 0.814 $\pm$ 0.008 0.246 $\pm$ 0.008 0.301 $\pm$ 0.002
@@
Table 14: Contrastive augmentation ablations (5 seeds). Per-view recomputation improves over stable SPD.
@@ 4.7 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact indices released) [14]. OGB and TDC use official splits [6, 7]. We remove any benchmark molecules from the pretraining corpus via scaffold match (Bemis-Murcko scaffolds [1] computed via RDKit [9] on the union of all train/val/test molecules across 11 benchmarks;  20k removed total, e.g.,  5k for MoleculeNet,  10k for OGB,  5k for TDC). Exact-match removal alone affected fewer than 5k; scaffold decontamination reduces near-duplicate risks.
@@ 5 Discussion
Tables 810 and 5 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 13 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage ¿95%), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 6). Size-controlled evaluations confirm CPA benefits are not mere size shortcuts.
Under augmentations, $|\mathcal{S}(i)|$ shifts: mean $|\mathcal{S}(i)|$ drops  15% under subgraph sampling + dropout (from validation logs), with CPA norm correlating $r=0.48$ post-augmentation (vs. 0.52 static), supporting dynamic adaptation.

### v4

@@ 4.5 Chemistry-Aware Augmentation Ablations
Attribute masking 0.818 $\pm$ 0.007 0.803 $\pm$ 0.008 +0.015 0.542 $\pm$ 0.020 0.590 $\pm$ 0.030 -0.048 0.920 $\pm$ 0.008
Valency-constrained 0.820 $\pm$ 0.007 0.804 $\pm$ 0.008 +0.016 0.540 $\pm$ 0.020 0.590 $\pm$ 0.030 -0.050 0.922 $\pm$ 0.008
@@ 4.7 Pretraining Corpus and Standardization
  • Source datasets: Random sample of 40M from ZINC20 clean-leads subset (approximately 220M SMILES from the full $\sim$1.4B-compound ZINC20 database) [8] and full ChEMBL35 (2.4M molecules) [25].
@@ 4.8 Optimization and Training Schedule
Per-view SPD recomputation (via multi-source BFS) increases pretraining cost by $\sim$15% $\pm$ 2% in wall-clock time; ablations under reduced budgets (10%, 25%, 100%) show CPA gains persist (Table 13).
@@ 4.9 Faithful Graphormer Baseline
For faithful comparison, we implement Graphormer with global attention ($K=\infty$), full SPD binning (up to max observed, clipped at 20, covering $99\%$ SPDs in our corpus), shortest-path edge encodings as learnable biases for each edge type along the shortest path (see appendix for the exact path-bias implementation), and centrality biases. Pretraining uses the same dual objectives (mask + contrast) and corpus. Hyperparameters match our backbone (12 layers, 512 dim, 8 heads). This isolates sparsity and CPA effects from bias design. See appendix for full implementation details. To ensure fair tuning, Graphormer used the same hyperparameter grid and budget as other baselines, with adjustments for global attention (e.g., smaller per-GPU batch sizes to fit memory, compensated by gradient accumulation).
@@ 5 Experiments
All transformer-family baselines (CardinalGraphFormer, SparseGraphormer-K3, GraphMAE-K3, GraphCL, and Graphormer) share the same backbone depth/width, pretraining corpus, optimization budget, and fine-tuning protocol unless explicitly noted. Differences are summarized in Table 5.
@@ 5.3 Efficiency
Table 6: Efficiency measured on A100 80GB GPUs (4 GPUs, DP=data parallel, per-GPU batch=1024, global batch=4096), bf16 (bfloat16 mixed precision), no activation checkpointing, FlashAttention-2 kernel. Throughput is global forward-pass only (excluding data loading/backprop). Peak memory is per-GPU CUDA allocated. Both settings use the same attention implementation; $K=3$ is implemented as an attention mask (not block-sparse kernels); compute savings therefore come primarily from shorter effective attention spans on larger graphs and reduced quadratic scaling in the padded attention length $L$ (Table 7). For typical drug-like molecules, $K=3$ is near-global (median coverage $95\%$); substantial sparsity benefits emerge on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering). Median coverage (%) is computed as $\mathrm{median}_{G}\!\left(\frac{1}{N}\sum_{i\in V}\frac{|\mathcal{S}(i)|}{N}\right)\times 100$.
Table 6 reports measured efficiency. For typical drug-like molecules (dominant in benchmarks), $K=3$ is primarily a regularizer, yielding near-global coverage (median $95\%$). Meaningful efficiency gains emerge on larger graphs ($\approx 15\%$ of the pretraining corpus after filtering), where sparsity provides practical compute savings and memory efficiency. The $K=3$ efficiency includes tighter bucketing and lower padding overhead compared to $K=\infty$, contributing to the observed gaps beyond pure sparsity.
@@ 5.4 Results
Statistical significance uses paired bootstrap (10,000 resamples per task/seed, with identical resample indices across models to preserve pairing, deltas averaged across seeds); denotes 95% CI excludes zero vs SparseGraphormer-K3 after Holm correction across the 11 benchmark tasks (detailed CIs and adjusted p-values in Appendix Table 16; all primary deltas remain significant post-correction).
@@
Task (metric) CardinalGraphFormer $\Delta$ SparseGraphormer-K3 GraphMAE-K3 GraphCL Graphormer D-MPNN GIN
ESOL (RMSE $\downarrow$) 0.542 $\pm$ 0.020 -0.056 0.598 $\pm$ 0.030 0.580 $\pm$ 0.020 0.612 $\pm$ 0.030 0.609 $\pm$ 0.032 0.550 $\pm$ 0.020 0.589 $\pm$ 0.030
Lipophilicity (RMSE $\downarrow$) 0.629 $\pm$ 0.010 -0.029 0.658 $\pm$ 0.020 0.662 $\pm$ 0.020 0.672 $\pm$ 0.020 0.668 $\pm$ 0.021 0.640 $\pm$ 0.010 0.650 $\pm$ 0.020
BBBP (ROC-AUC $\uparrow$) 0.938 $\pm$ 0.009 +0.020 0.918 $\pm$ 0.011 0.922 $\pm$ 0.010 0.910 $\pm$ 0.012 0.912 $\pm$ 0.011 0.930 $\pm$ 0.009 0.915 $\pm$ 0.011
Tox21 (ROC-AUC $\uparrow$) 0.921 $\pm$ 0.008 +0.013 0.908 $\pm$ 0.009 0.909 $\pm$ 0.007 0.902 $\pm$ 0.010 0.899 $\pm$ 0.009 0.915 $\pm$ 0.008 0.905 $\pm$ 0.009
ClinTox (ROC-AUC $\uparrow$) 0.954 $\pm$ 0.011 +0.015 0.939 $\pm$ 0.017 0.938 $\pm$ 0.015 0.932 $\pm$ 0.018 0.935 $\pm$ 0.017 0.952 $\pm$ 0.010 0.940 $\pm$ 0.016
@@
Table 8: MoleculeNet scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3 (see Appendix Table 16 for details). Additional baselines reproduced under our protocol.
@@
Dataset (metric) CardinalGraphFormer $\Delta$ SparseGraphormer-K3 GraphMAE-K3 GraphCL Graphormer D-MPNN GIN
ogbg-molhiv (ROC-AUC $\uparrow$) 0.819 $\pm$ 0.007 +0.017 0.802 $\pm$ 0.009 0.803 $\pm$ 0.008 0.798 $\pm$ 0.010 0.800 $\pm$ 0.010 0.810 $\pm$ 0.007 0.795 $\pm$ 0.009
ogbg-molpcba (AP $\uparrow$) 0.304 $\pm$ 0.002 +0.010 0.294 $\pm$ 0.002 0.295 $\pm$ 0.002 0.292 $\pm$ 0.003 0.300 $\pm$ 0.003 0.302 $\pm$ 0.002 0.290 $\pm$ 0.002
@@
Table 9: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks; see Appendix Table 16). Additional baselines reproduced under our protocol.
@@
Benchmark (metric) CardinalGraphFormer $\Delta$ SparseGraphormer-K3 GraphMAE-K3 GraphCL Graphormer D-MPNN GIN
Caco2_Wang (MAE $\downarrow$) 0.241 $\pm$ 0.007 -0.024 0.265 $\pm$ 0.009 0.268 $\pm$ 0.008 0.270 $\pm$ 0.009 0.272 $\pm$ 0.010 0.250 $\pm$ 0.007 0.260 $\pm$ 0.008
hERG (ROC-AUC $\uparrow$) 0.898 $\pm$ 0.004 +0.028 0.870 $\pm$ 0.005 0.872 $\pm$ 0.004 0.865 $\pm$ 0.006 0.862 $\pm$ 0.007 0.880 $\pm$ 0.004 0.868 $\pm$ 0.005
Clearance_Hepatocyte_AZ (Spearman $\uparrow$) 0.538 $\pm$ 0.017 +0.013 0.525 $\pm$ 0.020 0.518 $\pm$ 0.018 0.520 $\pm$ 0.021 0.518 $\pm$ 0.020 0.530 $\pm$ 0.017 0.522 $\pm$ 0.019
Clearance_Microsome_AZ (Spearman $\uparrow$) 0.621 $\pm$ 0.011 +0.007 0.614 $\pm$ 0.012 0.615 $\pm$ 0.012 0.610 $\pm$ 0.013 0.612 $\pm$ 0.012 0.618 $\pm$ 0.011 0.612 $\pm$ 0.012
@@
Table 10: TDC ADMET scaffold split results (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. For $\downarrow$ metrics, negative $\Delta$ indicates improvement. significant vs SparseGraphormer-K3 (see Appendix Table 16 for details). Additional baselines reproduced under our protocol.
@@ 5.5 Size-Controlled and Size-Shift Evaluations
ESOL (RMSE $\downarrow$) Small 0.548 $\pm$ 0.022 0.610 $\pm$ 0.030 0.855 $\pm$ 0.048
Mid 0.541 $\pm$ 0.020 0.600 $\pm$ 0.030 0.828 $\pm$ 0.052
Large 0.529 $\pm$ 0.021 0.585 $\pm$ 0.030 0.822 $\pm$ 0.050
Size-shift 0.539 $\pm$ 0.021 0.605 $\pm$ 0.030 0.835 $\pm$ 0.050
BBBP (AUC $\uparrow$) Small 0.935 $\pm$ 0.009 0.915 $\pm$ 0.011 0.650 $\pm$ 0.020
Mid 0.938 $\pm$ 0.009 0.916 $\pm$ 0.011 0.635 $\pm$ 0.020
Large 0.939 $\pm$ 0.009 0.919 $\pm$ 0.011 0.618 $\pm$ 0.020
Size-shift 0.937 $\pm$ 0.009 0.917 $\pm$ 0.011 0.634 $\pm$ 0.020
hERG (AUC $\uparrow$) Small 0.895 $\pm$ 0.004 0.865 $\pm$ 0.005 0.698 $\pm$ 0.012
Mid 0.897 $\pm$ 0.004 0.868 $\pm$ 0.005 0.687 $\pm$ 0.009
Large 0.900 $\pm$ 0.004 0.875 $\pm$ 0.005 0.682 $\pm$ 0.011
Size-shift 0.897 $\pm$ 0.004 0.869 $\pm$ 0.005 0.689 $\pm$ 0.010
Tox21 (AUC $\uparrow$) Small 0.918 $\pm$ 0.008 0.905 $\pm$ 0.009 0.750 $\pm$ 0.010
Mid 0.920 $\pm$ 0.008 0.907 $\pm$ 0.009 0.742 $\pm$ 0.010
Large 0.925 $\pm$ 0.008 0.910 $\pm$ 0.009 0.728 $\pm$ 0.010
Size-shift 0.921 $\pm$ 0.008 0.907 $\pm$ 0.009 0.740 $\pm$ 0.010
molpcba (AP $\uparrow$) Small 0.302 $\pm$ 0.002 0.292 $\pm$ 0.002 0.200 $\pm$ 0.010
Mid 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.192 $\pm$ 0.010
Large 0.306 $\pm$ 0.002 0.296 $\pm$ 0.002 0.178 $\pm$ 0.010
Size-shift 0.304 $\pm$ 0.002 0.294 $\pm$ 0.002 0.190 $\pm$ 0.010
@@ 5.6.1 Cardinality-Control Baselines
Per-view SPD (primary) 0.819 $\pm$ 0.007 0.241 $\pm$ 0.007 0.304 $\pm$ 0.002
Stable SPD (baseline) 0.814 $\pm$ 0.008 0.246 $\pm$ 0.008 0.301 $\pm$ 0.002
@@
Table 14: Contrastive augmentation ablations (5 seeds). Per-view recomputation improves over stable SPD and serves as the primary augmentation strategy.
@@ 5.7 Data, Splits, and Overlap Removal
MoleculeNet tasks use 80/10/10 Bemis–Murcko scaffold splits (scaffolds sorted by frequency, round-robin assignment, matching DeepChem protocol; exact split indices are included in the reproducibility package described in Code and Data Availability) [20]. OGB and TDC use official splits [6, 7]. We remove any benchmark molecules from the pretraining corpus via scaffold match (Bemis-Murcko scaffolds [1] computed via RDKit [11] on the union of all train/val/test molecules across 11 benchmarks;  20k removed total, e.g.,  5k for MoleculeNet,  10k for OGB,  5k for TDC). Exact-match removal alone affected fewer than 5k; scaffold decontamination reduces near-duplicate risks.
@@ 6 Discussion
Tables 810 show consistent improvements attributable to the CPA channel (objective-matched comparison to SparseGraphormer-K3) and dual pretraining (SparseGraphormer-K3 vs GraphMAE-K3). Table 13 indicates the CPA contribution is substantially larger than individual static centrality components, with degree-normalized CPA consistent with a dynamic cardinality interpretation. For typical drug-like molecules (median $N\approx 22$), $K=3$ covers nearly all nodes (median coverage $95\%$), suggesting gains arise primarily from structured biases and CPA rather than strict sparsity enforcement; on larger graphs ($\approx 15\%$, computed from the pretraining corpus after filtering), $K$ provides meaningful compute savings (Table 6). Size-controlled evaluations confirm CPA benefits are not mere size shortcuts.
Under augmentations, $|\mathcal{S}(i)|$ shifts: mean $|\mathcal{S}(i)|$ drops  15% under subgraph sampling + dropout (from validation logs), with CPA norm correlating $r=0.48\pm 0.04$ (95% CI: [0.40, 0.56]) post-augmentation (vs. $0.52\pm 0.03$ static), supporting dynamic adaptation.
@@
Per-Task Effect Sizes. Examining the per-task improvements (Table 810), hERG exhibits the largest CPA benefit (+0.028 AUC), while Clearance_Microsome_AZ shows the smallest (+0.007 Spearman). hERG, a molecular toxicity prediction task, may benefit disproportionately from cardinality signals as it involves discriminating between compounds with subtle structural differences in local neighborhoods; the dynamic support-size signal helps distinguish such molecular contexts. Conversely, clearance tasks, which depend on broader metabolic pathways, show diminishing returns from local cardinality information, suggesting the gain is task-dependent and correlated with the informativeness of local structural variation.
[…]
Future Work: Chemical Interpretability. A promising direction is interpreting the learned CPA gates and attention weights through the lens of chemical mechanisms, e.g., correlating high-weight neighbors with pharmacophoric features or reaction centers. Such analysis could provide actionable insights for medicinal chemists, bridging black-box neural predictions with human-interpretable chemical reasoning.